2. Consider the following model of the labour market. where w is the wage rate, Ld is labour demanded by the firms and Ls is labour supplied by workers What condition should δ satisfy in order for the second equation to be a reasonable labour supply function (i) What condition should satisfy in order for this system to have a unique equilibrium. (iii) Assume that δ = 1, express the systemin matrix form and use matrix algebra to find the equilibrium values of L and w.
Answer #1

i)

Labor supply should be an increasing function of wage. Thus the slope of L-w graph must be positive. Hence for a reasonable labor supply, δ>0.

ii)

In order for system to have a unique equilibrium -

L_d = L_s Rightarrow 2 - w^* = 1 + delta w^*

Rightarrow delta (w^*+1) = 1 Rightarrow delta = rac{1}{1+w^*} where w* is the equilibrium wage.

c)

If delta = 1, equations become -

L_d = 2 -w Rightarrow L_d + w = 2

L_s = 1 + w Rightarrow L_s - w = 1

In the equilibrium, Ld = Ls = L

Hence, these equations can be written in the matrix form as -

egin{bmatrix} 1 & 1 1& -1 end{bmatrix}egin{bmatrix} L w end{bmatrix}1

L + w1

Solving them, we get:

Adding both rows : 2L = 3 => L = 3/2

and thus w = 2 - 3/2 = 1/2